OpenStudy (anonymous):

working on practice... help! Solve the equation 7n - 2 = 5n + 10 for n. A. 1 B. 1.5 C. 4 D. 6

OpenStudy (uri):

7n - 2 = 5n + 10 7n-5n=10+2 Solve? @alicat28

OpenStudy (anonymous):

thank you so much! i am having an off day :)

OpenStudy (uri):

:)

OpenStudy (anonymous):

that aint right tho

OpenStudy (uri):

@Kittywuvschu It isn't completely solved.

OpenStudy (anonymous):

its 6

OpenStudy (uri):

yeah 6.

OpenStudy (anonymous):

@uri you are a lifesaver :)

OpenStudy (uri):

:D